Suppose that a monopolistically competitive restaurant is currently serving 250 meals per day (the output where MR = MC). At that output level, ATC per meal is $10 and consumers are willing to pay $13 per meal.

Instructions: Enter your answers as whole numbers.

a. What is the size of this firm’s profit or loss?

b. Will there be entry or exit?

      (Click to select)  Exit  Entry

     Will this restaurant’s demand curve shift left or right?

      (Click to select)  Right  Left

c. Suppose that the allocatively efficient output level in long-run equilibrium is 210 meals. In long-run equilibrium, suppose that this restaurant charges $11 per meal for 180 meals and that the marginal cost of the 180th meal is $9. What is the size of the firm’s economic profit?

d. Suppose that the allocatively efficient output level in long-run equilibrium is 210 meals. In long-run equilibrium, suppose that this restaurant charges $11 per meal for 180 meals and that the marginal cost of the 180th meal is $9. Is the deadweight loss for this firm greater than or less than $60?

      (Click to select)  Less than  Greater than
Answer #1

a) Profit = P-ATC*Q

=13-10*250

=750

b) Yes,there will be entry of new firms because the profits are positive

The restaurant's demand curve will shift to the left as a result.

c) The economic profit in the long run would be equal to zero because the firm operates where P=minimum of ATC

d) DWL = 0.5*base*height

=0.5*(210-180)*(11-9)

=0.5*30*2

=30

So it is less than $60